OpenStudy (anonymous):

Show that for all integers m and n, with m ≠ ±n, ∫_(-π)^π cos⁡(mθ)cos⁡(nθ)dθ=0. In English that is "the integral from negative pi to pi of..." If it helps, please note this is in the section involving integration by partial fractions and trig substitution.

OpenStudy (anonymous):

\[\cos p \ \cos q=\frac{1}{2} (\cos(p+q)+\cos(p-q))\]

OpenStudy (anonymous):

Is that just a trig identity that should be memorized? I do not see that in this section? What exactly am I supposed to do with that?

OpenStudy (turingtest):

just integrate what @mukushla showed you and evaluate

OpenStudy (turingtest):

you can also make use of the fact that cosine is an even function, and for even functions\[\int_{-a}^af(x)dx=2\int_0^af(x)dx\]to save some trouble in the evaluation

OpenStudy (anonymous):

Thanks! Both of you!